OpenStudy (anonymous):

divide and check 18x9 - 36 x^6 + 36 x^4 / 9x^4

OpenStudy (anonymous):

\[ \dfrac{18x^9-36x^6+36x^4}{9x^4}\\ \dfrac{9(2x^9-4x^6+4x^4)}{9x^4}\\ \dfrac{9x^4(2x^5-4x^2+4)}{9x^4}\\ 2x^5-4x^2+4 \]

OpenStudy (anonymous):

good answer nb